OpenStudy (anonymous):

A parabola has equation y - 7 = 3(x + 2)^2. Rewrite the equation in the form Ax^2 + Bxy + Cy^2 + Dx + Ey + F = 0.

OpenStudy (mertsj):

Expand the binomial (x+2)^2 , put all terms on the left and in the proper order. Use 0 for a coefficient if there is no such term.

OpenStudy (anonymous):

what about the -7?

OpenStudy (anonymous):

check me, please: 3x^2+12x+-1y+19=0

OpenStudy (mertsj):

Hang on. I will.

OpenStudy (mertsj):

y-7=3(x+2)^2= y-7=3(x^2+4x+4) y-7=3x^2+12x+12 3x^2+12x-y+19=0

OpenStudy (mertsj):

Yep. I agree.